A circular swimming pool has a diameter of 8 M and a depth of 2 m. what is the volume of the swimming pool??
AND HOW??
THANKS

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

Look at the swimming pool as a cylinder
diameter = 8 m ===> radius = 4m
depth/height = 2m
..
Volume V of cylinder = pi*r^2*h
V= pi*4^2*2
V= 16pi
V= 50.2 m^3
m.ananth@hotmail.ca

Answer by ikleyn(53427)

.
A circular swimming pool has a diameter of 8 M and a depth of 2 m. what is the volume of the swimming pool??
AND HOW??
THANKS
~~~~~~~~~~~~~~~~~~


        The solution, the calculations and the answer in the post by @manaqnth all are incorrect.
        I came to bring a correct and accurate solution.


Look at the swimming pool as a cylinder
diameter = 8 m ===> radius = 4m
depth/height = 2m
..
Volume V of cylinder = pi*r^2*h
V= pi*4^2*2
V= 32pi
V= 100.48 m^3

Solved correctly.